# Equivalence and Separation between Heard-Of and Asynchronous Message-Passing Models

### v1

@@ 7 Randomized protocols with a non-adaptive adversary
In this section, we extend our study to randomized protocols under a non-adaptive adversary, which fixes process crashes and message omissions independently of the protocol’s random choices. Randomization can break symmetry among processes, but it cannot repair lost communication: once two processes are silenced with respect to each other, their random choices evolve independently, and their outputs may diverge. We prove that the equivalence between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ for colorless tasks persists in this setting, while the separation for colored tasks remains. Thus, the expressive boundary between the two models is structural, determined by the flow of information rather than by deterministic constraints.
More specifically, we assume processes are equipped with their own local coins, which they may flip an arbitrary number of times during the local computation of each step. Together with an initial configuration and schedule, the sequences of local coin flips determine a run. An overview of randomized protocols can be found in [Aspnes03AsyncCons].
@@
We choose the same message schedule and input vector $(x_{1},\dots,x_{n-2},a,b)$ as in the proof of Theorem 14, with $a$ and $b$ to be chosen later. Let $R$ be the set of runs having the chosen input vector and schedule. In run $r\in R$, let $c^{\prime}(r)$ be the sequence of random choices of the processes $p_{1},\dots,p_{n-2}$ and let $c_{n-1}(r)$ and $c_{n}(r)$ be those of $p_{n-1}$ and $p_{n}$. Write $\mathcal{C}^{\prime}$ for the set of sequences of random choices $c^{\prime}(r)$. Once $x_{1},\dots,x_{n-2}$ are fixed, the new names of $p_{n-1}$ and $p_{n}$ depend on their own initial names, their own sequences of random choices and those of $p_{1},\dots,p_{n-2}$, and thus can be written as $\delta(a,c_{n-1}(r),c^{\prime}(r))$ and $\delta(b,c_{n}(r),c^{\prime}(r))$. We use the fact that all processes almost-surely decide distinct non-$\bot$ values and the law of total probability [Billingsley1995Prob, Theorem 34.4] to obtain
@@
Once $c^{\prime}(r)$ is fixed to $c^{\prime}$ and values for $a$ and $b$ are chosen, the new names $\delta(a,c_{n-1}(r),c^{\prime})$ and $\delta(b,c_{n}(r),c^{\prime})$ are independent random variables due to $c_{n-1}(r)$ and $c_{n}(r)$ being independent. Moreover they must be chosen from a set of k=N(n2)k=N-(n-2) names. By setting $c=1/2k$ in Lemma 17, there exists a choice of initial names $a$ and $b$ from a set of $L=L(N,c)$ names such that
@@ 8 Summary and Discussion
This work establishes a precise correspondence between the asynchronous message-passing model with crash failures, $\mathrm{AMP}_{f}$, and the round-based Heard-Of model with message omissions, $\mathrm{HO}_{f}$. The two coincide for colorless tasks when $n>2f$, and for colored tasks only when $f=1$ (and $n>2$). The distinction stems from silenced processes in $\mathrm{HO}_{f}$: when several processes become permanently unheard, their views diverge and they may reach incompatible decisions. With a single such process, compatibility can still be preserved; with more, it cannot. The results extend to randomized algorithms with non-adaptive adversaries, indicating that the limitations arise from communication structure rather than determinism.
@@ Appendix B Compactness of $\mathrm{HO}_{f}$
In this section, we prove that the set of runs of a protocol in $\mathrm{HO}_{f}$ with a finite number of possible initial configurations is compact with respect to the longest-common-prefix metric. Given two sequences $\alpha=(\alpha_{k})_{k\in\mathbb{N}_{0}}$ and $\beta=(\beta_{k})_{k\in\mathbb{N}_{0}}$, their longest-common-prefix distance [AS84] is:
@@
To prove compactness of the set of schedules, we use Tychonoff’s theorem (see, e.g. [Bourbaki:top, Chapter I, §9, no. 5, Theorem 3]).
@@ Appendix C $\mathrm{AMP}$ is equivalent to the FLP definition
In order to show that our definition of $\mathrm{AMP}_{f}$ is equivalent to the $\mathrm{FLP}_{f}$ definition, we show that each model can simulate the other. For the simulation of $\mathrm{AMP}_{f}$ in $\mathrm{FLP}_{f}$, the broadcasting of a message in $\mathrm{AMP}_{f}$ can be simulated in $\mathrm{FLP}_{f}$ by sending it individually to all processes, and a message $m$ is said to be received in $\mathrm{AMP}_{f}$ if $m$ is received in $\mathrm{FLP}_{f}$.
@@
1:When $\operatorname{amp-send}_{i}(m)$ occurs:
@@
For the simulation of $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$, the sending of a message $m$ in $\mathrm{FLP}_{f}$ to $p_{j}$ can be simulated in $\mathrm{AMP}_{f}$ by broadcasting the message $\left\langle m,j\right\rangle$, and a message $m$ is said to be received by process $p_{i}$ in $\mathrm{FLP}_{f}$ if $p_{i}$ receives $\left\langle m,i\right\rangle$ in $\mathrm{AMP}_{f}$.
@@
2:  Enable $\operatorname{amp-send}_{i}(\left\langle m,j\right\rangle)$

### v2

@@ 7 Randomized protocols with a non-adaptive adversary
In this section, we extend our study to randomized protocols under a non-adaptive adversary, which fixes process crashes and message omissions independently of the protocol’s random choices. Randomization can break symmetry among processes, but it cannot repair lost communication: once two processes are silenced with respect to each other, their random choices evolve independently and their outputs may diverge. We prove that the equivalence between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ for colorless tasks persists in this setting, while the separation for colored tasks remains. Thus, the expressive boundary between the two models is structural, determined by the flow of information rather than by deterministic constraints.
More specifically, we assume processes are equipped with their own local coins, which they may flip an arbitrary number of times during the local computation of each step. Together with an initial configuration and schedule, the sequences of local coin flips determine a run. An overview of randomized protocols can be found in [3].
@@
Let $N\geq 1$ be a positive integer and let 0<c<1/N0<c<{1}/{N}. There exists an integer $L=L(N,c)$ with the following property: for every collection $X_{1},\dots,X_{L}$ of $L$ pairwise independent random variables in the set $\{a_{1},\dots,a_{N}\}$, there exist distinct indices $a,b\in\{1,\dots,L\}$ such that $\mathbb{P}(X_{a}=X_{b})\geq c$.
@@
We choose the same message schedule and input vector $(x_{1},\dots,x_{n-2},a,b)$ as in the proof of Theorem 14, with $a$ and $b$ to be chosen later. Let $R$ be the set of runs having the chosen input vector and schedule. In run $r\in R$, let $c^{\prime}(r)$ be the sequence of random choices of the processes $p_{1},\dots,p_{n-2}$ and let $c_{n-1}(r)$ and $c_{n}(r)$ be those of $p_{n-1}$ and $p_{n}$. Write $\mathcal{C}^{\prime}$ for the set of sequences of random choices $c^{\prime}(r)$. Once $x_{1},\dots,x_{n-2}$ are fixed, the new names of $p_{n-1}$ and $p_{n}$ depend on their own initial names, their own sequences of random choices and those of $p_{1},\dots,p_{n-2}$, and thus can be written as $\delta(a,c_{n-1}(r),c^{\prime}(r))$ and $\delta(b,c_{n}(r),c^{\prime}(r))$. We use the fact that all processes almost-surely decide distinct non-$\bot$ values and the law of total probability [8, Theorem 34.4] to obtain
@@
Once $c^{\prime}(r)$ is fixed to $c^{\prime}$ and values for $a$ and $b$ are chosen, the new names $\delta(a,c_{n-1}(r),c^{\prime})$ and $\delta(b,c_{n}(r),c^{\prime})$ are independent random variables because of $c_{n-1}(r)$ and $c_{n}(r)$ being independent. Moreover they must be chosen from a set of k=N(n2)k=N-(n-2) names. By setting $c=1/2k$ in Lemma 17, there exists a choice of initial names $a$ and $b$ from a set of $L=L(N,c)$ names such that
@@ 8 Summary and Discussion
This work establishes a precise correspondence between $\mathrm{AMP}_{f}$, the asynchronous message-passing model with crash failures, and $\mathrm{HO}_{f}$, the round-based Heard-Of model with message omissions. The two coincide for colorless tasks when $n>2f$, and for colored tasks only when $f=1$ (and $n>2$). The distinction stems from silenced processes in $\mathrm{HO}_{f}$: when several processes become permanently unheard, their views diverge and they may reach incompatible decisions. With a single such process, compatibility can still be preserved; with more, it cannot. The results extend to randomized algorithms with non-adaptive adversaries, indicating that the limitations arise from communication structure rather than determinism.
@@
Acknowledgements
Dhrubajyoti Ghosh and Thomas Nowak are supported by the ANR projects DREAMY (ANR-21-CE48-0003) and COSTXPRESS (ANR-23-CE45-0013). Hagit Attiya is supported by the Israel Science Foundation (grant number 25/1849). Armando Castañeda is supported by research projects DGAPA-PAPIIT IN108723 and IN103126, and SECIHTI CBF-2025-I-393.
@@ Appendix B Compactness of $\mathrm{HO}_{f}$
In this section, we prove that the set of runs of a protocol in $\mathrm{HO}_{f}$ with a finite number of possible initial configurations is compact with respect to the longest-common-prefix metric. Given two sequences $\alpha=(\alpha_{k})_{k\in\mathbb{N}_{0}}$ and $\beta=(\beta_{k})_{k\in\mathbb{N}_{0}}$, their longest-common-prefix distance [2] is:
@@
To prove compactness of the set of schedules, we use Tychonoff’s theorem (see, e.g. [9, Chapter I, §9, no. 5, Theorem 3]).
@@ Appendix C $\mathrm{AMP}$ is equivalent to the FLP definition
In order to show that our definition of $\mathrm{AMP}_{f}$ is equivalent to the $\mathrm{FLP}_{f}$ definition, we construct simulations between the two models (see Section 2 for the definition of a simulation). For the simulation of $\mathrm{AMP}_{f}$ in $\mathrm{FLP}_{f}$, the broadcasting of a message in $\mathrm{AMP}_{f}$ can be simulated in $\mathrm{FLP}_{f}$ by sending it individually to all processes, and a message $m$ is said to be received in $\mathrm{AMP}_{f}$ if $m$ is received in $\mathrm{FLP}_{f}$.
@@
1:When $\operatorname{amp-bc}_{i}(m)$ occurs:
@@
For the simulation of $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$, the sending of a message $m$ in $\mathrm{FLP}_{f}$ by $p_{i}$ to $p_{j}$ can be simulated in $\mathrm{AMP}_{f}$ by having $P_{i}$ broadcasting the message $\left\langle m,j\right\rangle$, and a message $m$ is said to be received by process $p_{i}$ in $\mathrm{FLP}_{f}$ if $P_{i}$ receives $\left\langle m,i\right\rangle$ in $\mathrm{AMP}_{f}$.
@@
2:  Enable $\operatorname{amp-bc}_{i}(\left\langle m,j\right\rangle)$
@@
Note that in a simulated run $\alpha$ of the system described by Algorithm 5, the last message $m$ sent by a faulty process $p_{i}$ to a non-faulty process $p_{j}$ is not necessary received by $p_{j}$, preventing $\alpha$ from being valid in $\mathrm{FLP}_{f}$. This is because of the Faulty Quasi-Liveness property of $\mathrm{AMP}_{f}$ whereby $P_{i}$’s last broadcast $\left\langle m,j\right\rangle$ is not necessarily received by all machines, in particular $P_{j}$. Fortunately this does not prevent non-faulty processes in $\alpha$ from deciding, since for them, $\alpha$ is indistinguishable from a valid $\mathrm{FLP}_{f}$ run $\alpha^{\prime}$ that is almost identical to $\alpha$ except that faulty simulated processes crash before they can do the incomplete send, e.g., $\operatorname{flp-send}_{i}(m,j)$ in the case of $p_{i}$.